OpenStudy (anonymous):

solve by completing the square 3x^2-5x-10=0

OpenStudy (anonymous):

\[3x^2-5x=10\] \[x^2-\frac{5}{3}x=\frac{10}{3}\] \[(x-\frac{5}{6})^2=\frac{10}{3}+\frac{25}{36}=\frac{145}{36}\] \[x-\frac{5}{6}=\pm\frac{\sqrt{145}}{6}\]

OpenStudy (anonymous):

.....this is one messy problem